Jamie spent 1/4 of her money on 20 notebooks and 48 pens. The cost of each notebook was 3 times the cost of the pen. She bought some more notebooks with 2/3 of the remaining money. How many notebooks did Jamie buy altogether?
Answer by ikleyn(52835)

.

Jamie bought 20 notebooks and 48 pens. 1 notebook cost 3 times the cost of the pen.


Hence 20 notebooks and 48 pens are equivalent to 20 + 48/3 = 20 + 16 = 36 notebooks.


Thus, you can think that she spent 1/4 of her money for 36 notebooks.


After she spent 1/4 of her money, she left with 3/4 of her money.


2/3 of 3/4 is  1/2.


So, she additionally spent 1/2 of her money for additional notebooks.


Hence, the number of additional notebooks is 2 times 36, i.e. 72.


Thus, in all, she bought 20 + 72 = 92 notebooks, altogether.      ANSWER

Solved.